OpenStudy (lisa123):

The symbol {x} represents the sum of all the integers from 1 to x. For example {4}=1+2+3+4=10 If {a}-{b}=8, What could b equal?

OpenStudy (lisa123):

Please help me @Loser66

OpenStudy (loser66):

{a}-{b}={8} {a}-8={b} hence a must be 8 and b =7 . let check a =8 {8}= 1+2+3+4+5+6+7+8 {8}-8 = 1+2+3+4+5+6+7 ={7}={b}

OpenStudy (loser66):

the first line is {a}-{b}=8, not ={8}

OpenStudy (loser66):

However, I don't know how to put it in logic yet. Just use "try and error" method to find it out. It might have other solution but I didn't find it out yet.

OpenStudy (loser66):

@myko What do you think?

OpenStudy (lisa123):

This is an SAT question and the options are 8 7 6 2 1

OpenStudy (loser66):

oh, it is quite easy now. So, 7 is the solution

OpenStudy (anonymous):

i would say: \({a}-{b}= |n_1 +n_2 \cdots a|=8\) where n are consecutive integers. So a + this integers must sum 8. Now just check which can match this.

OpenStudy (anonymous):

i ment {a}-{b} in the first line

OpenStudy (lisa123):

If a=8 and b=7 Then is it 36-28=8?

OpenStudy (anonymous):

yes, thats the only possible solution

OpenStudy (lisa123):

Okay it makes sense now thank you gym for helping me :)

OpenStudy (anonymous):



OpenStudy (lisa123):

There are none @myko

OpenStudy (anonymous):

there are. Just 8 alone